OpenStudy (anonymous):

Write in slope intercept form: (2, -1), m= 2/3 I know the slope is "m" (so, 2/3x), but I'm having problems with the process of figuring out "b". I have: -1 = 2/3(2) + b -1 = 4/6 + b

OpenStudy (saifoo.khan):

You are going the correct way.. can u solve further?

OpenStudy (unklerhaukus):

just isolate b and sub it in

OpenStudy (unklerhaukus):

your almost there

OpenStudy (anonymous):

That's where I get stuck, on every problem. I think: -1/6 = -4/6 + b -2/6 = b -1/3 = b It seems wrong, somehow...

OpenStudy (precal):